OpenStudy (anonymous):

Write the equation y - 2 = 2(x - 3) in standard form. A. y - 2x = 4 B. 2x - y = -6 C. 2x + y = 6 D. 2x - y = 4

OpenStudy (anonymous):

d

OpenStudy (mathstudent55):

Distribute the 2 on the right side. Then collect like terms. y - 2 = 2(x - 3) y - 2 = 2x - 6 2x - y = 4